An elevator of mass 1500 kg is at rest and then begins to ascend at an acceleration of 1.0 m / s² . A constant frictional force of 800 N retards its motion upward . ( a ) Find the tension in the cable carrying the elevator . ( b ) Find the minimum power delivered by the motor at t = 4s while the elevator is moving upward . ( c ) What will be the reading on a bathroom scale if a 50 kg person stands on it while the elevator is ascending ?

1
Expert's answer
2022-03-17T13:55:13-0400

Answer

a) "ma=T-mg"

"T=m(g+a) \\=1500*10.8\\=16200N"

b) power

P=Et

c) F=ma